OpenStudy (anonymous):

In august Amy spent 2/5 of the money in her checking account. In september she spent 1/3 of the remainder. if she has $400 left, what was in her account originally? A) 100 B) 1000 C) 1200 D) 1600 E) 200

OpenStudy (across):

Let's denote Amy's initial balance as: \[B\]In August, Amy spent 2/5 of her money. That means she has 3/5 of her original balance: \[B-(2/5)B=(3/5)B\]In September, Amy spent 1/3 of the remainder. That means she spent 1/3 of the 3/5 remaining: \[(3/5)B-(1/3)((3/5)B)=(3/5)B-(1/5)B=(2/5)B\]She now has $400 left, this means that: \[(2/5)B=400\]Now solve for B and you obtain: \[B=$1000\]

OpenStudy (anonymous):

thank you very much across ^.^